OpenStudy (anonymous):

How do I find the radius of a quarter circle? The perimeter is 14.28.

OpenStudy (anonymous):

r≈2.27

OpenStudy (anonymous):

the radius is 2.27 ,

OpenStudy (anonymous):

How did you do it?

OpenStudy (anonymous):

Or what is the formula? I could figure it out from there.

OpenStudy (anonymous):

\[2\pi r\]

OpenStudy (anonymous):

 \[2 r+\frac{ 2 \pi r }{ 4 }= 14.28\]

OpenStudy (anonymous):

Cool, thnx.